Calorimetry can also be used to determine nutritional Calories, which are actually kilocalories. One (1) nutritional Calorie = 1000 calories, and 1 calorie = 4.184 J.

When 1.0 g of a snack cake is combusted in a constant-volume bomb calorimeter with a heat capacity (including its water) of 10.0 kJ/°C, the temperature of the calorimeter rises by 1.88°C. How many nutritional calories (Calories) are in one serving (28 g) of this snack cake?

Enter your answer to three significant figures in units of Calories.
Answer #1

Energy absorbed by the calorimeter = C * T = 10.0 * 1.88 = 18.8 KJ = 18.8 /4.184 * 103 cal = 4.4933 kcal

= 4.4933 nutritional cal

Energy in 28 g of cake = 28 * energy per 1 g of cake = 28 * 4.4933 = 125.8124 Nutritional calories